OpenStudy (anonymous):

A football is kicked toward the goal. The height of the ball is modeled by the function h(x) = -16t2 + 64t where t equals the time in seconds and h(x) represents the height of the ball at time t seconds. What is the axis of symmetry and what does it represent?

OpenStudy (anonymous):

x = 2; it takes 2 seconds to reach the maximum height and 2 seconds to fall back to the ground x = 2; it takes 2 seconds to reach the maximum height and 4 seconds to fall back to the ground x = 4; it takes 4 seconds to reach the maximum height and 4 seconds to fall back to the ground x = 4; it takes 4 seconds to reach the maximum height and 8 seconds to fall back to the ground

OpenStudy (johnweldon1993):

We can just find the vertex of the parabola this graph makes... \[\large -\frac{b}{2a} = -\frac{64}{2(-16)} = \frac{-64}{-32} = ?\] Now what does that represent?